How many 2-digit numbers less than 60 are even?
Answer by Edwin McCravy(20054)

How many 2-digit numbers less than 60 are even?
I'll show you two ways.

The way you do this depends on whether you are studying
arithmetic sequences or not.  If you are not, then do
it this way:

The smallest is 20 and the largest is 58

Those numbers go like this: 

               20,22,24,...,54,56,58

The numbers that are halves of those numbers go like this

               10,11,12,...,27,28,29

The numbers that are 9 less than those numbers go like this

               1 , 2, 3,...,18,19,20

Answer: 20.

-------------------

If you are taking algebra and studying arithmetic sequences, then
do it this way:

an = a1 + (n-1)d

an = 58, a1 = 20, d=2

58 = 20 + (n-1)(2)

38 = 2(n-1)

Divide both sides by 2

19 = n-1

20 = n

Answer = 20.

Edwin